OpenStudy (anonymous):

Find numerically y(2) for the differential equation y'=x+y^2, y(0)=0 by using the Euler method with h=1

OpenStudy (anonymous):

define a line at each point, using the slope at that point ...

OpenStudy (anonymous):

y' = 0 + 0^2 at 0,0 .. this is the line y=0, step to x=1 at x=1, y=0, y' = 1+0^2 = 1, y=x-1 ... at x=2, y=1 solve for the point (2,1) etc ...

OpenStudy (anonymous):

excel could prolly help with the iterations

OpenStudy (anonymous):

oh, lol it says to get to y(2) which i spose i got to with my example already :/

OpenStudy (anonymous):